[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Sun Jun 04, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Mon Jun 5 04:37:37 UTC 2017


HSU Shift Log for Sun Jun 04 2017 13:11:00 GMT-0400 (EDT)

Shift: H

Forecaster: M Brennan

Technical Issues Summary:

ATCF issues:

N-AWIPS issues:

AWIPS-2/NCP issues:

Other Techincal issues: There was GOES-16 outage from 16Z until 20Z today.  
Issue with loose tiles on the HSU ops floor near the printer is worsening  
and needs to be addressed.

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170605/9e9122ce/attachment-0001.html 


More information about the Ncep.nhc.nco_contacts mailing list